# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix D Welfare formulas used in Section 3
$\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}(g,z)&=-\log\left(\mathbb{E}\left[e^{-\psi_{I,0}^{\prime}p(h)-\widehat{\psi}_{I}(G,Z_{N})^{\prime}(X-p(H))}\big{|}\sigma(G,Z_{N})\right]\right)\big{|}_{(G,Z_{N})=(g,z)},\\ &=\psi_{I,0}^{\prime}p(h)-\frac{1}{2}\widehat{\psi}_{I}(g,z)^{\prime}(P_{I}+P_{X})^{-1}\widehat{\psi}_{I}(g,z)\\ &\qquad\qquad+\widehat{\psi}_{I}(g,z)^{\prime}(P_{I}+P_{X})^{-1}\left(P_{X}\mu_{X}+P_{I}g-(P_{I}+P_{X})p(h)\right),\end{split}$
where we used (65) and $X\sim N(\mu_{X},P_{X}^{-1})$. Therefore, (62) implies
(77) $\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}(g,z)&=\psi_{I,0}^{\prime}p(h)+\frac{1}{2}\left\|(P_{I}+P_{X})^{-1/2}\widehat{\psi}_{I}(g,z)\right\|^{2}.\end{split}$
The uninformed has the same signal type as the insider, giving
(78) $\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0}(h)&=\psi_{U,0}^{\prime}p(h)+\frac{1}{2}\left\|(P_{U}+P_{X})^{-1/2}\widehat{\psi}_{U}(h)\right\|^{2},\end{split}$
where $\widehat{\psi}_{U}$ is from (62). In the price impact equilibrium (recall (51) implies $\sigma(G,H_{\iota})=\sigma(G,Z_{N})$) we first obtain
@@
$\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0,\iota}(g,z)=-\log\left(\mathbb{E}\left[e^{-\psi_{I,0}^{\prime}p_{\iota}\left(\widehat{\psi}_{I},Z_{N}\right)-\widehat{\psi}_{I,\iota}^{\prime}\left(X-p_{\iota}\left(\widehat{\psi}_{I},Z_{N}\right)\right)}\big{|}\sigma(G,Z_{N})\right]\right)\big{|}_{(G,Z_{N})=(g,z)}.$
Using (6) and (65), we see that $-(1/\alpha_{I})\textrm{CE}^{I}_{0,\iota}(g,z)$ is (66), evaluated at $\widehat{\psi}_{I,\iota}$. Therefore
@@
(79) $\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0,\iota}(g,z)&=-\psi_{I,0}^{\prime}M\left(\psi_{I,0}-\frac{z}{\alpha_{I}}\right)+\psi_{I,0}^{\prime}V\\ &\qquad+\frac{1}{2}\left\|\left(M+M^{\prime}+(P_{I}+P_{X})^{-1}\right)^{1/2}\widehat{\psi}_{I,\iota}(g,z)\right\|^{2},\end{split}$
which below will be further specified plugging in for $M$ from (48) and $\widehat{\psi}_{I,\iota}$ from (67). For the uninformed agent $U$, as she is a price taker, the formulas are the same as in the non-impact case except with $H_{\iota},P_{U,\iota},p_{\iota}$ replacing $H,P_{U},p$. As such
(80) $\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0,\iota}(h_{\iota})&=\psi_{U,0}^{\prime}p_{\iota}(h_{\iota})+\frac{1}{2}\left\|\left(P_{U,\iota}+P_{X}\right)^{-1/2}\widehat{\psi}_{U,\iota}(h_{\iota})\right\|^{2},\end{split}$
where $\widehat{\psi}_{U,\iota}$ is from (53). Using the above general formulas, the first Lemma identifies interim welfare in the price taking equilibrium under Assumption A.4 and using the notation in (68).
Lemma D.1.
$\begin{split}&\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}(g,z)=\psi_{I,0}^{\prime}\mu_{X}+\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}P_{X}^{1/2}(g-\mu_{X})+\frac{R}{1-R}P_{X}^{-1/2}\frac{z}{\alpha_{I}}\\ &\qquad\qquad-\frac{R(1+\beta)}{(1-R)(\beta+\lambda_{I})}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad+\frac{\lambda_{U}^{2}R(1-R)^{2}}{2(\lambda_{U}R+\beta+\lambda_{I})^{2}}\bigg{\|}P_{X}^{1/2}(g-\mu_{X})-\frac{\beta+\lambda_{I}}{\lambda_{U}(1-R)}P_{X}^{-1/2}\frac{z}{\alpha_{I}}+\frac{1+\beta}{\lambda_{U}(1-R)}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
And
$\begin{split}&\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0}(h)=\psi_{U,0}^{\prime}\mu_{X}+\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}\psi^{\prime}_{U,0}P_{X}^{-1/2}\bigg{(}P_{X}^{1/2}(h-\mu_{X})-\frac{R(1+\beta)}{(1-R)(\beta+\lambda_{I})}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad+\frac{R(1+\beta)\lambda_{I}^{2}(1-R)^{2}}{2(R+\beta)(\lambda_{U}R+\beta+\lambda_{I})^{2}}\bigg{\|}P_{X}^{1/2}(h-\mu_{X})-\frac{R+\beta}{\lambda_{I}(1-R)}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Proof of Lemma D.1.
Proposition C.1 and (55) imply $p(h)=\mu_{X}-P_{X}^{-1}\widehat{\Pi}+m_{g}\left(h-\mu_{X}+P_{X}^{-1}\widehat{\Pi}\right)$, where $m_{g}$ is from (32). Therefore, from (60) and (62) we find
$\begin{split}\widehat{\psi}_{I}(g,h)&=P_{X}\mu_{X}+P_{I}g-(P_{I}+P_{X})\left(\mu_{X}-P_{X}^{-1}\widehat{\Pi}+m_{g}\left(g+P_{I}^{-1}\frac{z}{\alpha_{I}}-\mu_{X}+P_{X}^{-1}\widehat{\Pi}\right)\right).\end{split}$
Grouping terms by $g-\mu_{X}$, $z/\alpha_{I}$ and $\widehat{\Pi}$ we obtain
$\begin{split}g-\mu_{X}&:\quad P_{I}-(P_{I}+P_{X})m_{g}=\frac{\lambda_{U}(1-R)}{\lambda_{U}R+\beta+\lambda_{I}}P_{X},\\ \frac{z}{\alpha_{I}}&:\quad-(P_{I}+P_{X})m_{g}P_{I}^{-1}=-\frac{\beta+\lambda_{I}}{\lambda_{U}R+\beta+\lambda_{I}}1_{d},\\ \widehat{\Pi}&:\quad(P_{I}+P_{X})P_{X}^{-1}-(P_{I}+P_{X})m_{g}P_{X}^{-1}=\frac{1+\beta}{\lambda_{U}R+\beta+\lambda_{I}}1_{d}.\end{split}$
Therefore
@@
(81) $\begin{split}\widehat{\psi}_{I}(g,z)&=\frac{1}{\lambda_{U}R+\beta+\lambda_{I}}\left(\lambda_{U}(1-R)P_{X}\left(g-\mu_{X}\right)-(\beta+\lambda_{I})\frac{z}{\alpha_{I}}+(1+\beta)\widehat{\Pi}\right).\end{split}$
Similar calculations show
(82) $\begin{split}p(h)&=\mu_{X}+\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}\bigg{(}g-\mu_{X}+\frac{R}{1-R}P_{X}^{-1}\frac{z}{\alpha_{I}}-\frac{R(1+\beta)}{(1-R)(\beta+\lambda_{I})}P_{X}^{-1}\widehat{\Pi}\bigg{)}.\end{split}$
As $(P_{I}+P_{X})^{-1}=RP_{X}^{-1}$ we obtain from (77)
$\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}(g,z)&=\psi_{I,0}^{\prime}\mu_{X}+\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}\psi_{I,0}^{\prime}\bigg{(}g-\mu_{X}+\frac{R}{1-R}P_{X}^{-1}\frac{z}{\alpha_{I}}-\frac{R(1+\beta)}{(1-R)(\beta+\lambda_{I})}P_{X}^{-1}\widehat{\Pi}\bigg{)}\\ &\ +\frac{R\lambda_{U}^{2}(1-R)^{2}}{2(\lambda_{U}R+\beta+\lambda_{I})^{2}}\bigg{\|}P_{X}^{1/2}(g-\mu_{X})-\frac{\beta+\lambda_{I}}{\lambda_{U}(1-R)}P_{X}^{-1/2}\frac{z}{\alpha_{I}}+\frac{1+\beta}{\lambda_{U}(1-R)}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
The result follows factoring out $P_{X}^{-1/2}$ from the linear expression. The calculations are similar for the uniformed trader. From (60) and (62) we obtain
$\begin{split}\widehat{\psi}_{U}(h)&=P_{X}\mu_{X}+P_{U}h-(P_{U}+P_{X})\left(\mu_{X}-P_{X}^{-1}\widehat{\Pi}+m_{g}\left(h-\mu_{X}+P_{X}^{-1}\widehat{\Pi}\right)\right).\end{split}$
From (72) we see $P_{U}+P_{X}=(R+\beta)/(R(1+\beta))P_{X}$, and hence grouping terms by $h-\mu_{X}$ and $\widehat{\Pi}$ we obtain
$\begin{split}h-\mu_{X}&:\quad P_{U}-(P_{U}+P_{X})m_{g}=-\frac{\lambda_{I}(1-R)}{\lambda_{U}R+\beta+\lambda_{I}}P_{X},\\ \widehat{\Pi}&:\quad(1-m_{g})\left(P_{U}+P_{X}\right)P_{X}^{-1}=\frac{R+\beta}{\lambda_{U}R+\beta+\lambda_{I}}1_{d}.\end{split}$
Therefore,
$\widehat{\psi}_{U}(h)=-\frac{\lambda_{I}(1-R)}{\lambda_{U}R+\beta+\lambda_{I}}P_{X}(h-\mu_{X})+\frac{R+\beta}{\lambda_{U}R+\beta+\lambda_{I}}\widehat{\Pi}.$
As $(P_{U}+P_{X})^{-1}=R(1+\beta)/(R+\beta)P_{X}^{-1}$, we conclude from (30), (78) and (82)
$\begin{split}&\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0}(h)=\psi_{U,0}^{\prime}\mu_{X}+\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}\psi_{U,0}^{\prime}\left(h-\mu_{X}-\frac{R(1+\beta)}{(1-R)(\beta+\lambda_{I})}P^{-1}_{X}\widehat{\Pi}\right)\\ &\qquad+\frac{R(1+\beta)\lambda^{2}_{I}(1-R)^{2}}{2(R+\beta)(\lambda_{U}R+\beta+\lambda_{I})^{2}}\bigg{\|}P_{X}^{1/2}(h-\mu_{X})-\frac{R+\beta}{\lambda_{I}(1-R)}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
The result follows factoring out $P_{X}^{-1/2}$ from the linear expression.
The next Lemma identifies interim welfare in the price impact equilibrium.
Lemma D.2.
@@
$\begin{split}&\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0,\iota}(g,z)=\psi_{I,0}^{\prime}\mu_{X}+\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\widehat{y}RP_{X}^{-1/2}\frac{z}{\alpha_{I}}-\frac{(\widehat{y}+1)^{2}+\beta}{\lambda_{U}(\widehat{y}+1)^{2}}P_{X}^{-1/2}\widehat{\Pi}\\ &\qquad\qquad-\frac{\beta\widehat{y}}{(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{)}\\ &\qquad+\frac{(1-R)^{2}}{2R(1+2\widehat{y})}\bigg{\|}P_{X}^{1/2}(g-\mu_{X})-\frac{\widehat{y}R}{1-R}P_{X}^{-1/2}\frac{z}{\alpha_{I}}+\frac{(\widehat{y}+1)^{2}+\beta}{\lambda_{U}(1-R)(\widehat{y}+1)^{2}}P_{X}^{-1/2}\widehat{\Pi}\\ &\qquad\qquad+\frac{\widehat{y}\left(R(\widehat{y}+1)^{2}+\beta\right)}{(1-R)(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{\|}^{2}.\end{split}$
And
$\begin{split}&\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0,\iota}(h_{\iota})=\psi_{U,0}^{\prime}\mu_{X}+\psi_{U,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{(1-R)\widehat{y}}{1+2\widehat{y}}P_{X}^{1/2}(h_{\iota}-\mu_{X})\\ &\quad+\frac{\widehat{y}\lambda_{I}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+2\widehat{y})(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}-\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\quad+\frac{1}{2}\frac{\lambda_{I}^{2}(1-R)^{2}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}R(1+2\widehat{y})^{2}(\beta+R(1+\widehat{y})^{2})}\bigg{\|}P_{X}^{1/2}(h_{\iota}-\mu_{X})+\frac{(\beta+R(1+\widehat{y})^{2})\widehat{y}}{(1+\widehat{y})^{2}(1-R)}P_{X}^{-1/2}\psi_{I,0}\\ &\quad-\frac{\beta+R(1+\widehat{y})^{2}}{(1+\widehat{y})(1-R)\lambda_{I}}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Proof of Lemma D.2.
We start by analyzing $\widehat{\psi}_{I,\iota}$ from (67), using (48) and (59). This gives
$\widehat{\psi}_{I,\iota}(g,z)=\frac{1}{1+2\widehat{y}}\left(P_{I}g+P_{X}\mu_{X}-(P_{I}+P_{X})V+2\widehat{y}\left(\psi_{I,0}-\frac{z}{2\alpha_{I}}\right)\right).$
Using the notation in (68), plugging in for $V$ from (71), and recalling (55) we obtain
$\begin{split}\widehat{\psi}_{I,\iota}(g,z)&=\frac{1}{1+2\widehat{y}}\bigg{(}\frac{1-R}{R}P_{X}(g-\mu_{X})+\frac{1}{R}\widehat{\Pi}-\frac{\beta+\lambda_{I}(1+\widehat{y})^{2}}{\lambda_{U}R(1+\widehat{y})^{2}}(\psi_{I,0}-\widehat{\Pi})\\ &\qquad\qquad+2\widehat{y}\left(\psi_{I,0}-\frac{z}{2\alpha_{I}}\right)\bigg{)}.\end{split}$
Grouping terms by $P_{X}(g-\mu_{X}),\widehat{\Pi}$ and $z/\alpha_{I}$ we obtain
$\begin{split}P_{X}(g-\mu_{X})&:\frac{1-R}{R(1+2\widehat{y})},\qquad\widehat{\Pi}:\frac{(1+\widehat{y})^{2}+\beta}{\lambda_{U}R(1+\widehat{y})^{2}(1+2\widehat{y})},\qquad\frac{z}{\alpha_{I}}:-\frac{\widehat{y}}{1+2\widehat{y}}.\end{split}$
The $\psi_{I,0}$ terms are
$\frac{1}{R(1+2\widehat{y})}\left(2R\widehat{y}-\frac{\beta+\lambda_{I}(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}\right).$
As $\widehat{y}$ solves (69) we know
(83) $\frac{\beta+\lambda_{I}(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}=R\widehat{y}-\frac{\beta\widehat{y}}{(1+\widehat{y})^{2}}.$
Plugging this in and simplifying gives the $\psi_{I,0}$ terms
$\frac{\widehat{y}(\beta+R(1+\widehat{y})^{2})}{R(1+2\widehat{y})(1+\widehat{y})^{2}}.$
Therefore
(84) $\begin{split}\widehat{\psi}_{I,\iota}(g,z)&=\frac{1-R}{R(1+2\widehat{y})}P_{X}(g-\mu_{X})+\frac{(1+\widehat{y})^{2}+\beta}{\lambda_{U}R(1+\widehat{y})^{2}(1+2\widehat{y})}\widehat{\Pi}-\frac{\widehat{y}}{1+2\widehat{y}}\frac{z}{\alpha_{I}}\\ &\qquad\qquad+\frac{\widehat{y}(\beta+R(1+\widehat{y})^{2})}{R(1+2\widehat{y})(1+\widehat{y})^{2}}\psi_{I,0}.\end{split}$
Next, specifying $M,P_{I},P_{X}$ from (59) and $V$ from (71) in (79) gives

### v2

@@ B.4. Utility formulas used in Section 3
(86) $\begin{split}\frac{\mathbb{P}\left[X\in dx\big{|}\sigma(G,Z_{N})\right]}{\mathbb{P}\left[X\in dx\right]}&=\left(\frac{e^{-\frac{1}{2}x^{\prime}P_{I}x+x^{\prime}P_{I}g}}{\mathbb{E}\left[e^{-\frac{1}{2}X^{\prime}P_{I}X+X^{\prime}P_{I}g}\right]}\right)\bigg{|}_{g=G},\\ \frac{\mathbb{P}\left[X\in dx\big{|}\sigma(H_{k})\right]}{\mathbb{P}\left[X\in dx\right]}&=\left(\frac{e^{-\frac{1}{2}x^{\prime}P_{U,k}x+x^{\prime}P_{U,k}h}}{\mathbb{E}\left[e^{-\frac{1}{2}X^{\prime}P_{U,k}X+X^{\prime}P_{U,k}h}\right]}\right)\bigg{|}_{h=H_{k}}.\end{split}$
This allows us to define the “certainty equivalents”, valid under Assumption A.4, and using the notation in (77)
@@
(87) $\begin{split}\textrm{CE}^{o}_{I,0}(G)&:=-\alpha_{I}\log\left(\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(G,Z_{N})\right]\right)=-\frac{\alpha_{I}R}{2}\left(\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}-2\widehat{\Pi}^{\prime}\left(\mu_{X}+\frac{1-R}{R}G\right)\right),\\ \textrm{CE}^{o}_{U,\iota}(H_{\iota})&:=-\alpha_{U}\log\left(\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(H_{\iota})\right]\right)=-\frac{\alpha_{U}}{2(1+p_{U,\iota})}\left(\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}-2\widehat{\Pi}^{\prime}(\mu_{X}+p_{U,\iota}H_{\iota})\right),\\ \textrm{CE}^{o}_{U}(H)&:=-\alpha_{U}\log\left(\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(H)\right]\right)=-\frac{\alpha_{U}}{2(1+p_{U})}\left(\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}-2\widehat{\Pi}^{\prime}(\mu_{X}+p_{U}H)\right),\end{split}$
where $p_{U,\iota},p_{U}$ are from (79). Lastly, extending (26) to the general setup of Appendix B.2, we define the certainty equivalent absent private information
@@
(88) $\textrm{CE}^{i}_{nsn}:=\alpha_{i}\log\left(\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\right]\right)=\alpha_{i}\left(\widehat{\Pi}^{\prime}\mu_{X}-\frac{1}{2}\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}\right),\qquad i\in\left\{I,U\right\}.$
Price impact certainty equivalents
In this section our goal is to prove the following.
Proposition B.1.
Using the notation of (77), and $\widehat{y}$ from Proposition 1.1, the interim price impact certainty equivalents of Section 3 are
@@
(89) $\begin{split}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})&=\textrm{CE}^{o}_{I,0}(G)+\frac{\alpha_{I}R}{2(1+2\widehat{y})}\left|\frac{1-R}{R}P_{X}^{1/2}(G-p_{{\mathbb{Q}}_{0}})-\widehat{y}P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2},\\ \textrm{CE}^{U}_{0,\iota}(H_{\iota})&=\textrm{CE}^{o}_{U,0}(H_{\iota})+\frac{\alpha_{U}\lambda_{I}^{2}(1-R)^{2}(\beta+(1+\widehat{y})^{2})}{2\lambda_{U}^{2}R(\beta+R(1+\widehat{y})^{2})(1+2\widehat{y})^{2}}\left|P_{X}^{1/2}(H_{\iota}-p_{{\mathbb{Q}}_{0}})\right|^{2}.\end{split}$
The ex-ante price impact certainty equivalents of Section 3 are
@@
(90) $\begin{split}\textrm{CE}^{I}_{0-,\iota}&=\textrm{CE}^{I}_{nsn}+\frac{\alpha_{I}}{2}\log\left(1+\frac{(1-R)(\beta/R+\widehat{y}^{2})}{(1+2\widehat{y})\beta}\right),\\ \textrm{CE}^{U}_{0-,\iota}&=\textrm{CE}^{U}_{nsn}+\frac{\alpha_{U}}{2}\log\left(1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2}}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\right).\end{split}$
Proof of Proposition B.1.
We start with the insider. From (41), (43), (47) and (48)